How to find a turning point if you are given this expression : -2x^2+40x-600

Answer by stanbon(75887)

How to find a turning point if you are given this expression : -2x^2+40x-600
----
The "turning point" or vertex occurs when x = -b/(2a)
----
Your Problem::
x = -b/(2a) = -40/(2*-2) = 10
Then y = f(10) = -2(10)^2 + 40(10) - 600 = -200+400-600 = -400
---
Ans: Vertex is (10,-400)
--------
Cheers,
Stan H.
--------------